howw does:
tan ([pi] + alpha)= tan (alpha)
i tried using the sum and difference formula like it says in the directions but it isnt helping much.
i use the formula:
tan ([pi]) + tan(alpha) all divided by 1- tan( [pi] tan alpha)

Answer by lwsshak3(11628)

howw does:
tan ([pi] + alpha)= tan (alpha)
i tried using the sum and difference formula like it says in the directions but it isnt helping much.
i use the formula:
tan ([pi]) + tan(alpha) all divided by 1- tan( [pi] tan alpha)
**
using x for alpha
tan(x+π)=(tanx+tanπ)/(1-tanxtanπ)
tanπ=0
tan(x+π)=(tanx+0)/(1-tanx*0)=tanx/1=tanx
tan(x+π)=tanx
..
Physically, in a unit circle,you are advancing the angle π radians which places the angle x in a quadrant where the reference angle is equal to x and the sign of the tan function does not change. For example, if x were 21º, advancing it 180º places it in quadrant III where tan>0, and the reference angle is 21º